# Topological Strings in SU(3) Gauge Theory at Finite Temperature

### v1

@@ IV Numerical Setup and Results
For all the patches, there is a steep rise in $\sigma/T^{2}$ vs $\beta$ near $\beta_{c}$ and a linear rise for large $\beta$. The rise in $\sigma/T^{2}$ is higher for larger patches. For smaller patches, the contribution of the string core is expected to dominate the free energy. On the other hand, for large patches, the contribution from the domain walls dominates as they can extend all the way to the boundary of the system. All these results suggest that the free energy contribution of the core region rises slower in $\beta$ compared to the contribution from the domain walls. Moreover, for larger patches, the string tension and interface tension results agree with an overall $\beta$ independent scale factor, suggesting the dominance of the interface tension. This also implies that string configurations with larger winding numbers will have higher string tension due to the presence of more interfaces. In addition to this, there will be higher contribution from the string core for higher winding numbers.
@@ V Conclusion
In this work, we have conducted a detailed first-principles non-perturbative study of string configurations in the deconfined phase of $SU(3)$ gauge theory, originating from the spontaneous breaking of $Z_{3}$ symmetry. Using the Polyakov loop effective potential model, we argued that these strings are topologically stable. In the lattice simulations, our focus is mainly on the winding one string, which results from the junction of three domain walls. We used specific boundary conditions that lead to the formation of a string with three domain walls connected. Our methodology, based on indirect free energy estimation through action differences, was validated by recovering known results for the $Z_{3}$ interface tension, in good agreement with previous studies. This suggests that our approach reliably captures the essential physics of the string configurations. Subsequently, we calculated the string tension, $\sigma/T^{2}$ as a function of $\beta$. Our results show that the string tension rises steeply near $\beta_{c}$ and linearly for large $\beta$. The $Z_{3}$ string being a global defect, the string tension rises with the size of the area of the cross section. For a $U(1)$ global string, the tension approaches $log(r)$, where $r$ is the radius of the cross section. In this present case, the string tension for larger radius is dominated by the interface tension, hence rises linearly in r.

### v2

@@ IV.2 $N_{\tau}=4$ results
For both $N_{\tau}=2$ and $N_{\tau}=4$, the rise in $\sigma/T^{2}$ is larger for bigger radial patches. For smaller patches the contribution of the string core dominates the free energy, while for larger patches the contribution from the domain walls becomes dominant as they extend to the boundary of the system. These observations indicate that the free energy contribution from the core grows more slowly with $\beta$ compared to the contribution from the domain walls. Moreover for larger patches the string tension and interface tension results agree up to an overall $\beta$ independent scale factor, confirming that the interface tension dominates the string tension.
@@ V Conclusion
In this work, we have conducted a detailed first-principles non-perturbative study of string configurations in the deconfined phase of $SU(3)$ gauge theory, originating from the spontaneous breaking of $Z_{3}$ symmetry. Using the Polyakov loop effective potential model, we argued that these strings are topologically stable. In the lattice simulations, our focus is mainly on the winding one string, which results from the junction of three domain walls. We used specific boundary conditions that lead to the formation of a string with three domain walls connected. Our methodology, based on indirect free energy estimation through action differences, was validated by recovering known results for the $Z_{3}$ interface tension, in qualitatively good agreement with previous studies. This suggests that our approach reliably captures the essential physics of the string configurations. Subsequently, we calculated the string tension, $\sigma/T^{2}$ as a function of $\beta$. Our results show that the string tension rises steeply near $\beta_{c}$. For large $\beta$(weak coupling limit) the string tension for $N_{\tau}=2$ and $N_{\tau}=4$ approaches $\sqrt{\beta}$ for large patches as the domain walls dominate over the string core. The $Z_{3}$ string being a global defect, the string tension rises with the size of the area of the cross section. For a $U(1)$ global string, the tension approaches $log(r)$, where $r$ is the radius of the cross section. In this present case, the string tension for larger radius is dominated by the interface tension, hence rises linearly in r.